what is the resistance of a 30 m silver wire with a 2 mm diameter


1
Expert's answer
2022-03-09T13:30:01-0500

Length of silver wire = 30 m

Diameter of silver wire is = 2mm =2×10-3

area of cross section of wire = "\pi r^2= 3.14×(2×10^{-3})^2"

A=12.56×10-6

Resistivity of silver wire = 1.59×10-8


"R=\frac{\rho l}{A}"


"R=\frac{1.59×10^{-8}× 30}{12.56×10^{-8}}"


R= 3.797 "\Omega"